Hi! I have a question that I'm having trouble with. I started to work on it but I'm not too sure if I'm supposed to take the perpendicular bisector or not. Here's the question:
A chord "AB" is 9cm from the center of the circle. The circle has a radius of 15cm. What is the length of "AB"?
If you could help me out, that would be great!
Thanks!
Ryan Desmond

Answer by stanbon(75887)

A chord "AB" is 9cm from the center of the circle. The circle has a radius of 15cm. What is the length of "AB"?
--------------
Draw the picture including the line segment to the chord.
That line segment is the perdicular bisector of the chord.
Call it's length "x".
Draw radii from the center to A and from the center to B.
You have a right triangle with hyp=15 and perp bisec = 9
The 3rd side = sqrt(15^2-9^2)=sqrt(225-81)=sqrt(144)=12
That 3rd side is half the length of the chord.
The length of the chord is 24.
Cheers,
Stan H.